OpenStudy (anonymous):

slope intercept form of -9x+10y=-9

OpenStudy (sleepyjess):

Welcome to OpenStudy! How do you think you would start this?

OpenStudy (anonymous):

@jasminemarieee ?

OpenStudy (anonymous):

Slope intercept form is: y = mx + b -9x + 10y = -9 So we want to get 'y' by itself on one side of the equation, so let's start by adding 9x to both sides. Can you do that? @jasminemarieee

OpenStudy (bradely):

Slope intercept is y = mx+c Add 9x both sides -9x+10y+9x=9x-9 10y=9x-9 Divide each side by 10 y=(9/10)x-(9/10) Step by step solution http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

oops! i meant add -9x :P

OpenStudy (anonymous):

Actually @DuckDynsatyfan923 we have to add 9x, not subtract.

OpenStudy (anonymous):

:P brain fart

OpenStudy (sleepyjess):

Subtracting a negative number is the same as adding though, subract -9x = +-9x

OpenStudy (anonymous):

That's still wrong..lol

OpenStudy (anonymous):

That was pointing to @DuckDynastyfan923 's comment, not yours @sleepyjess lol

OpenStudy (anonymous):

For some reason my tags aren't working..I have to type the whole tag out..

OpenStudy (sleepyjess):

That also happens to me

OpenStudy (anonymous):

@jasminemarieee are you still here?

OpenStudy (anonymous):

Anyway, adding -9x is the same thing as subtracting..

OpenStudy (anonymous):

goodness im sorry im here!

OpenStudy (anonymous):

No problem, Slope intercept form is: y = mx + b -9x + 10y = -9 So we want to get 'y' by itself on one side of the equation, so let's start by adding 9x to both sides. Can you do that? @jasminemarieee

OpenStudy (anonymous):

@iGreen. okay how would i do that? there is an x and it confuses me

OpenStudy (anonymous):

-9x + 10y = -9 add -9x on both sides to get 10y = 9x -9

OpenStudy (anonymous):

We just add it like anything else, but we don't combine it with anything: -9x + 10y = -9 Add 9x to both sides: 10y = -9 + 9x See? We just added it to -9.

OpenStudy (anonymous):

ohhhhhh that makes sense!!

OpenStudy (anonymous):

then divide both sides by 10. can you do that?

OpenStudy (anonymous):